Evaluate the following integrals using residue theorem: (a) ∫∞−∞ x² /(x² +9)(x² +4)²dx

1
Expert's answer
2022-05-31T11:12:57-0400

"\int_{-\infin}^{\infin}\cfrac{x^2dx}{(x^2+9)(x^2+4)^2}"

Let "F(x) = \cfrac{x^2}{(x^2+9)(x^2+4)^2}"

"F \in C(\mathbb{R})" because "\forall x \in \mathbb{R} \ x^2+9 \neq 0" and "(x^2+4)^2\neq 0"

Consider "F(z) = \cfrac{z^2}{(z^2+9)(z^2+4)^2}, z \in \mathbb{C}"

Find poles of "F(z):"

"z = \pm 3i" - simple pole

"z = \pm2i" - order 2 pole

So, using residue theorem:

"\int_{-\infin}^{\infin}\cfrac{x^2dx}{(x^2+9)(x^2+4)^2} = 2\pi i(\mathop{\mathrm{Res}}_{z = 2i}F(z)+\mathop{\mathrm{Res}}_{z = 3i}F(z))"

Find the residuals:

"\mathop{\mathrm{Res}}_{z = 3i}F(z) = \lim_{z \to3i} [F(z)(z-3i)] = \lim_{z \to3i} [\cfrac{z^2}{(z+3i)(z^2+4)^2}] = \cfrac{-9}{6i \cdot(-9+4)} = \cfrac{-3}{50i} = \cfrac{3i}{50}"


"\mathop{\mathrm{Res}}_{z = 2i}F(z) = \lim_{z \to 2i}[(\cfrac{z^2}{(z^2+9)(z+2i)^2})']"

Find derivative:

"(\cfrac{z^2}{(z^2+9)(z+2i)^2})' = \cfrac{2z(z^2+9)(z+2i)^2-z^2(2z(z+2i)^2+2(z+2i)(z^2+9))}{(z^2+9)^2(z+2i)^4}"

Then

"\mathop{\mathrm{Res}}_{z = 2i}F(z) = \lim_{z \to 2i}[\cfrac{2z(z^2+9)(z+2i)^2-z^2(2z(z+2i)^2+2(z+2i)(z^2+9))}{(z^2+9)^2(z+2i)^4}] =\\
=\cfrac{4i(-4+9)(4i)^2+4(4i(4i)^2+2(4i)(-4+9))}{(-4+9)^2(4i)^4} = \cfrac{-320i-96i}{6400} = \cfrac{-3i}{200}"

And finally:

"\int_{-\infin}^{\infin}\cfrac{x^2dx}{(x^2+9)(x^2+4)^2} = 2\pi i(\mathop{\mathrm{Res}}_{z = 2i}F(z)+\mathop{\mathrm{Res}}_{z = 3i}F(z)) = 2\pi i(\cfrac{3i}{50}-\cfrac{13i}{200}) = \cfrac{\pi}{100}"